OpenStudy (anonymous):

really need help on this operations with fractions the answer is there i just need step by step work they got the answer 5/16 + 1 3/4 + 4 1/2 +3/8 = 6 15/16

OpenStudy (anonymous):

Okay... since the answer is 6 15/16. We know that the denominator is going to be 16. To add the fractions you have to have the same denominator. How can we do that?

OpenStudy (anonymous):

do you know how do i have to get the same denominator

OpenStudy (anonymous):

Okay so since the first fraction's denominator is already a 16 you can leave that alone and focus on the others. What number can you use to multiply them and make the denominator 16? For the 1 3/4 and 4 1/2 , you first need to turn them into improper fractions.

OpenStudy (anonymous):

I'll do the 3/8 to show you have you would multiply to get the denominator 16. \[\frac{ 3 }{ 8 } \times 2 = \frac{ 6 }{ 16 }\]

OpenStudy (anonymous):

I'll do the 3/8 to show you how you would multiply to get the denominator 16.*

OpenStudy (anonymous):

i think i get it so you know the 1 3/4 i just take 3/4 and times it with 4 and its gives me 14/16 am I right

OpenStudy (anonymous):

No, you have to make it improper first so you would have to multiply the the denominator with the whole number and then add it with your numerator. Whatever you get as the answer will be the numerator over the original denominator 4.

OpenStudy (anonymous):

After that you can go ahead and multiply to get your new fraction. :)

OpenStudy (anonymous):

thx you can you help me on one other question this one can find a coomon number 7 11/4 + 1/6 + 3/4

OpenStudy (anonymous):

Write out your multiplications 2x4=8, 2x6=12, 3x4=12.

OpenStudy (anonymous):

ok thx i understand now

OpenStudy (anonymous):

No problem. :)